OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job?

OpenStudy (anonymous):

1247.44

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

or the you could get the area: (103*109)/(3*3) = 11227/9 = 1247.4

OpenStudy (anonymous):

vw